# Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable advances in supporting long-context comprehension and processing tasks. However, scaling the generation inference of LLMs to such long contexts incurs significant additional computation load, and demands a substantial GPU memory footprint to maintain the key-value (KV) cache of transformer-based LLMs. Existing KV cache compression methods, such as quantization, face memory bottlenecks as context length increases, while static-sized caches, such as selective eviction, suffer from inefficient policies. These limitations restrict deployment on consumer-grade devices like a single Nvidia 4090 GPU. To overcome this, we propose Locret, an efficient framework for long-context LLM inference that introduces retaining heads to evaluate the causal importance of KV cache units, allowing for more accurate eviction within a fixed cache size. Locret is fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets. During inference, we evict low-importance cache units along with a chunked prefill pattern, significantly reducing peak GPU memory usage. We conduct an extensive empirical study to evaluate Locret, where the experimental results show that Locret outperforms the recent popular and competitive approaches, including InfLLM, Quantization, SirLLM, and MInference, in terms of memory efficiency and the quality of generated contents — Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio compared to the full KV cache for Phi-3-mini-128K and Llama-3.1-8B-instruct. Additionally, Locret can be combined with other efficient inference methods, such as quantization and token merging. To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little additional system optimizations. The experimental code and the tool kit of Locret is provided at https://github.com/huangyuxiang03/Locret
@@ 1 Introduction
Large language models (LLMs) have revolutionized AI development and deployment (Zhao et al., 2023; Minaee et al., 2024). Recent advancements in LLMs’ ability to handle long-context tasks have further unlocked the potential of generative AI. State-of-the-art LLMs now support significantly extended context lengths, with GPT-4o (OpenAI, 2024) and Llama-3.1 (Dubey et al., 2024) handling 128K tokens, Yi (Young et al., 2024) and Claude-3 (Anthropic, 2024) supporting 200K tokens, and Gemini-1.5 (Reid et al., 2024) reaching 10 million tokens. These advances enable LLMs to tackle complex tasks like multi-hop reasoning (Li et al., 2024; Schnitzler et al., 2024), solving needle-in-a-haystack problems (Guerreiro et al., 2023; Wang et al., 2024a), and powering advanced LLM agents (Qin et al., 2023; Wang et al., 2024b) and AI-driven operating systems (Mei et al., 2024). However, deploying generative inference under long-context settings on consumer-grade GPUs requires innovative algorithmic and system optimizations to handle this new paradigm efficiently.
Compared to traditional short-context LLM inference, long-context LLM inference shifts the computing paradigm in two key ways: i) increased computational overhead for attention mechanisms: as context length grows, the computation required for obtaining attention scores increases quadratically, which results in a higher ratio of the computational budget in a transformer block; ii) higher memory footprint for key-value (KV) caching: longer contexts require larger KV caches, which dramatically increases the peak memory usage. These shifts demand innovative techniques to mitigate computational costs and manage memory usage effectively for long-context LLM inference.
Although various efforts have been made to overcome the bottleneck in LLM inference, these approaches fail to enable long-context inference on consumer-grade GPUs. Models with compact architectures, e.g. MiniCPM-128K (Hu et al., 2024b) and Phi-3-mini-128K (Abdin et al., 2024), reduce the computational load and memory usage, but cannot alleviate the KV cache burden in long-context scenarios. Similarly, techniques like LLM model weight quantization (Frantar et al., 2023; Lin et al., 2024; Ma et al., 2024), activation quantization (Dettmers et al., 2022; Xiao et al., 2023; Zhang et al., 2024c), or sparsification (Liu et al., 2023; Zhang et al., 2022) also fall short in effectively reducing the memory usage to a level supported by consumer-grade GPUs.
Recently, some specific optimizations have been proposed for long-context LLM inference. For example, sparse attention mechanisms (Jiang et al., 2024a; Ge et al., 2024; Lou et al., 2024) attempt to reduce runtime memory through conduct selected few calculation, and KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024) reduces cache size by applying low-bit storage. These methods can only offer limited compression rate, of which the core issue is that the KV cache grows linearly with sequence length, and the above methods do not adequately address this problem. On the contrary, combining chunked prefill with token-dropping techniques (Xiao et al., 2024b; Yang et al., 2024) could offer a more effective solution, as it maintains a static-sized cache where the memory usage can be bounded. However, current token-dropping and cache eviction methods (Zhang et al., 2024e; Liu et al., 2024a; Yao et al., 2024), whose token importance manually designed according to the inner statistics, suffer from accuracy loss and performance degradation due to inaccuracies in estimating token importance — The weakening correlation between local and global importance as sequences grow exacerbates this issue. To address these limitations, we propose a lightweight training-based paradigm that provides more accurate token importance scoring to tackle the long-context LLM inference problem. We highlight our contributions below:
Contribution 1: We propose a lightweight training-based paradigm for selective KV cache eviction for long-context LLM inference, with an offline training cost $<$1 GPU hours. We tackle the problem of KV cache eviction by a learning based approach. We introduce the retaining heads, with a small number of additional parameters, fine-tuned on top of the frozen backbone LLM using a minimal amount of data from standard long-context SFT datasets to estimate the causal importance of each cache unit. Such a training paradigm is able to provide accurate token importance scoring prediction and can be integrated with other efficient inference algorithms, e.g., quantization and token merging.
Contribution 2: We provide an efficient inference system implementation for Locret. We integrate the retaining head mechanism into a chunked prefill inference framework, where we maintain a static-size cache set through evicting cache units with low predicted importance to limit the GPU memory usage. Locret is able to preserve the most important cache units with the trained retaining heads, enabling precise attention approximation without compromising the inference latency. Locret is also applicable to all transformer-based LLMs and various hardware, as it requires minimal modifications to the model’s inference process and only utilizes dense operators.
Contribution 3: We conduct an extensive evaluation of Locret, which illustrates that Locret can not only obtain a comparable performance but also maintain inference efficiency. Locret achieves over a 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratio for Phi-3-mini-128K and Llama-3.1-8B-instruct, enabling full comprehension of long contexts on consumer-grade devices. To the best of our knowledge, Locret is the first framework capable of deploying Llama-3.1-8B or similar models on a single Nvidia 4090 GPU, enabling 128K long-context inference without compromising generation quality, and requiring little extra system optimizations.
@@ 2 Related Work
Efforts in long-context LLM inference can be categorized by algorithm and system optimizations:
@@
Algorithm optimizations. Optimizations aimed at reducing the size of the KV cache can generally be classified into three categories: quantization-based methods, token dropping methods, and sparsity-based methods. Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024; Zhang et al., 2024a), which store the KV cache in low-bit representations, require hardware support for these formats and may slow down inference due to the overhead of dequantization. Token dropping methods typically follow two main strategies: eviction or the use of an attention pool. Eviction-based approaches, such as H2(Zhang et al., 2024e), ScissorHands (Liu et al., 2024a), and SirLLM (Yao et al., 2024), rank tokens by certain statistical metrics to identify the most influential ones, discarding others to reduce memory usage. Attention pool-based methods (Nawrot et al., 2024; Rajput et al., 2024), such as StreamingLLM (Xiao et al., 2024b) and LoCoCo (Cai et al., 2024), compress multiple adjacent KV cache units into a single unit using a specially designed transformation. Sparsity-based methods (Ge et al., 2024; Jiang et al., 2024a; Yang et al., 2024; Lou et al., 2024; Lv et al., 2024) focus on leveraging the sparsity patterns of attention heads to reduce both computation and I/O. The combination of these approaches can be further enhanced by identifying specific attention patterns for each head and layer (Ge et al., 2024). For surveys of these methods, please refer to (Yuan et al., 2024; Kang et al., 2024; Shi et al., 2024).
@@
System optimizations. The challenge of long-context inference can also be alleviated from a system-level perspective. Offloading-based methods (Sheng et al., 2023; Xiao et al., 2024a; Wu et al., 2024; Lee et al., 2024) store the KV cache in CPU memory, retrieving only the most relevant chunks to the GPU before computing a new chunk. This approach reduces peak GPU memory usage, though at the cost of slower inference. Hardware-aware algorithms, such as flash attention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and page attention (Kwon et al., 2023), exploit GPU architecture (Ghorpade et al., 2012) to enable more efficient runtime memory management. In addition, reimplementing inference infra-structure in a more efficient programming language (llama.cpp, ; llama2.c, ; rustformers, ), or adopting disaggregated inference (Jiang et al., 2024b; Zhong et al., 2024; Qin et al., 2024; Hu et al., 2024a), can greatly enhance inference efficiency. Algorithmic optimizations can be seamlessly integrated into such systems (Agrawal et al., 2023; Lee et al., 2024). For instance, KTransformers (KVCache.AI, 2024) adopts the chunked offloading technique from InfLLM (Xiao et al., 2024a). However, system optimizations primarily focus on extending context length by leveraging hardware resources, rather than directly reducing the size of the KV cache.
3 Locret
@@ 3.1 Preliminaries
Transformer architecture. We define the model inference of transformer-based LLMs as follows. Given a token sequence $t_{1},t_{2},\cdots,t_{n}$, we denote the output hidden state of layer $i$ as $\textbf{H}^{(i)}$ and $\textbf{H}^{(0)}$ is the embeddings. Each transformer layer consists of an attention layer and an MLP layer. We assume the model follows a grouped-query attention (GQA) architecture (Ainslie et al., 2023), with $h$ query heads and a group size of $g$. For multi-head attention (MHA), $g$ is set to 1. The attention score of layer $i$’s head $j$ is calculated by $\mathbf{A}_{j}^{(i)}=\text{softmax}\left(\mathbf{Q}_{j}^{(i)}{\mathbf{K}_{\lceil{j/g}\rceil}^{(i)T}}/\sqrt{d_{m}}\right)\cdot\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}$, where $d_{m}$ represents the hidden size for each head and $\left[\mathbf{Q}_{j}^{(i)},\mathbf{K}_{j}^{(i)},\mathbf{V}_{j}^{(i)}\right]=\mathbf{H}^{(i-1)}\cdot\left[{\mathbf{W}_{j}^{Q(i)}},{\mathbf{W}_{j}^{K(i)}},{\mathbf{W}_{j}^{V(i)}}\right]$. Next, we compute $\mathbf{A}^{(i)}=\left[\mathbf{A}_{1}^{(i)},\cdots,\mathbf{A}_{h}^{(i)}\right]\cdot\mathbf{W}^{O(i)}$, finally followed by $\mathbf{H}^{(i)}=\text{MLP}(\mathbf{A}^{(i)})$.
@@
KV cache and chunked prefill. During the prefill stage, all prompt tokens are processed in a single forward pass, where $\mathbf{Q}^{(i)}$, $\mathbf{K}^{(i)}$, and $\mathbf{V}^{(i)}$ each have a sequence length of $n$. In the decoding stage, only a single token is processed across layers, utilizing the KV cache units to reduce computation. Chunked prefill is a method for reducing peak memory consumption by processing tokens in chunks over multiple passes, with the assistance of the KV cache. Taking both KV cache and chunked prefill into account, the attention calculation can be modified as Equation 1, where $B$ represents the number of tokens processed in a single model pass. For decoding, $B=1$, while for chunked prefill, $B$ corresponds to the chunk size. We denote the attention output for tokens $n+1,n+2,\cdots,n+B$ as $\mathbf{A}[n+1\colon n+B]$, and the attention output for the $k$-th token as $\mathbf{A}[k]$.
$\mathbf{A}{[n+1\colon n+B]}_{j}^{(i)}=\text{softmax}\left(\frac{\mathbf{Q}{[n+1\colon n+B]}_{j}^{(i)}{\mathbf{K}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)T}}}{\sqrt{d_{m}}}\right)\cdot\mathbf{V}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)}.$ (1)
@@
Cache eviction. Cache eviction in long-context inference is defined as follows. Here, we slightly abuse the notation of heads and layers, and treat the key-value vector pair of a single token within one head as the smallest cache unit. We denote the cache unit for the $k$-th token as $c_{k}=(\mathbf{K}[k],\mathbf{V}[k])$. Assume a memory budget $b$, representing the maximum number of cache units that can be stored in GPU memory at any given time. The abstract form of attention can then be written as $c_{k}=f(c_{1},c_{2},\cdots,c_{k-1})$. With limited cache capacity, this calculation can only be approximated by $\tilde{c}_{k}=f(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b^{\prime}\leq b$, and $p_{1},p_{2}\cdots,p_{b^{\prime}}\in\{1,2,\cdots,k-1\}$. Intuitively, the number of prior cache units involved cannot exceed the memory budget. When $b^{\prime}=b$, indicating the cache is full, one cache unit must be evicted. We select the unit to be evicted using a policy $p_{v}=\text{Policy}(\tilde{c}_{p_{1}},\cdots,\tilde{c}_{p_{b}};\tilde{c}_{k})$. In such stated problem, the key challenge of cache eviction is to develop an effective policy function that minimizes the approximation error $\|\tilde{c}_{k}-c_{k}\|$.
Figure 1: The framework of Locret. “$\mathbf{R}$” represents the retaining head. $\text{P}_{i}$ and $\text{A}_{i}$ correspond to the $i$-th prompt token and answer token. “t” represents the time step in chunked prefill, “$b$” represents the budget size, and “$n_{s}$” represents the length of the stabilizers.
3.2 Locret  Framework

### v2

@@ Abstract
Scaling the input context length of a large language model (LLM) incurs a significant increase in computation cost and memory footprint to maintain the attention key-value (KV) cache. Existing KV cache compression methods suffer from inefficient compression strategies and limited memory reduction effects, making it difficult for LLMs to conduct long-context inference on consumer-grade devices, especially when inferring long-context stream input. Such obstacles prevent consumer-grade devices from supporting more complex applications, creating challenges for the democratization of LLMs. To overcome this, we propose Locret, the first framework to create an eviction policy compatible with chunked prefill. By evaluating the causal importance of KV cache units by learnable retaining heads, Locret enables precise eviction of cache units, facilitating efficient long-context inference. In our extensive empirical studies, Locret outperforms the recent popular and competitive approaches in terms of memory efficiency and generation quality — Locret achieves up to 20×20$\times$20 × of KV cache compression ratio within less than $10\%$ performance loss. Furthermore, Locret achieves 128K+ long-context inference on a single NVIDIA 4090 GPU without compromising generation quality and only costs $<1$ GPU hour of additional training. We provide the code of Locret at https://github.com/huangyuxiang03/Locret.
Long-context Inference, Memory Efficient Inference, Large Language Models
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized generative AI (Zhao et al., 2023; Minaee et al., 2024), and the advancements of LLMs in handling long-context tasks have further unlocked the potential of generative AI. As a result, the context lengths supported by state-of-the-art LLMs have been significantly extended, such as GPT-4o (OpenAI, 2024) handling 128K tokens, Claude-3 (Anthropic, 2024) supporting 200K tokens, and Gemini-1.5 (Reid et al., 2024) even reaching 10M tokens. These improvements enable LLMs to tackle complex applications with extremely long or streaming inputs, such as multi-hop reasoning (Li et al., 2024b; Schnitzler et al., 2024), LLM-driven agents (Qin et al., 2024b; Wang et al., 2024), and AI-powered operating systems (Mei et al., 2024). Some recent efforts (Hu et al., 2024b; Abdin et al., 2024) have successfully deployed LLMs on consumer-grade end-side devices instead of cloud servers and conducted inference with limited context. We envision that unleashing the potential of long-context inference on consumer-grade devices will revolutionize the development of personalized AI applications and the democratization of LLMs. However, conducting long-context LLM inference on consumer-grade devices remains a challenging problem that requires algorithmic innovations and systematic optimizations.
Figure 1: For each prefix length of the context, this figure shows the consistency in evaluating the token importance of the prefix based on the full context and based on only the prefix without subsequent tokens. The consistency is defined as the intersection of the top 10% tokens of two evaluation methods divided by the number of top 10% tokens in the prefix. More details are in Appendix B.
As context length scales, the challenge of long-context LLM inference arises from two major aspects: the increased computational cost of the attention mechanism and the higher memory footprint due to the key-value (KV) cache. This leads to the failure of traditional optimizations targeting model backbones to provide a sufficient solution. Specifically, backbone-targeted optimizations, such as compacting model architectures (Hu et al., 2024b; Abdin et al., 2024) and quantizing model weights (Frantar et al., 2023; Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2024), fail to improve the efficiency of attention patterns or the KV cache, as attention’s quadratic complexity with respect to sequence length remains unaddressed. To this end, recent efforts have focused on optimizing attention patterns and the KV cache to achieve efficient long-context LLM inference.
Recent attention-targeting optimizations, including sparse attention (Jiang et al., 2024a; Ge et al., 2024; Lou et al., 2024) and KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024), show promising results to accelerate attention computation and reduce memory footprint. However, they fail to fundamentally address the core challenge: the KV cache grows linearly with context length. Layer-wise chunked prefill combined with attention sparsity (e.g., SnapKV (Li et al., 2024a)) can alleviate this problem to a certain extent. This technique typically performs cache eviction after the precise attention computation for each layer and requires access to the entire sequence. It can theoretically support longer sequences by limiting the maximum memory usage to a single layer’s KV cache, but it cannot handle streaming input whose length grows continually. The combination of cache eviction methods (Xiao et al., 2024c; Yang et al., 2024) and chunked prefill offers a more promising approach by maintaining a static cache size and supporting streaming input. Yet, as shown in Figure 1, existing eviction techniques like H2O (Zhang et al., 2024d) and SnapKV show significant discrepancies between local importance estimation and global importance estimation, i.e. it is hard to estimate the importance of each token only based its previous tokens. Instead, these methods require a large number of subsequent tokens to make an accurate estimation. Other methods like SirLLM (Yao et al., 2024) show great local-global estimation consistency but suffer from performance degradation.
To overcome these limitations, we propose a lightweight training-based paradigm, Locret, that provides a more accurate token importance estimation to select the victim during KV cache eviction, enabling efficient and scalable long-context LLM inference:
Contribution 1: We propose Locret, a lightweight training-based paradigm for selective KV cache eviction in long-context LLM inference. It introduces learnable retaining heads to estimate the causal importance score (CIS) for token selection, with an offline training cost of $<$1 GPU hour. Additionally, we present Locret-Q, a query-aware variant of Locret, slightly modified to handle query-driven tasks (e.g., long document question answering).
Contribution 2: We provide an efficient inference system implementation for Locret, integrating retaining heads into a chunked prefill framework. This integration limits GPU memory usage by evicting low-CIS KV cache units during the prefill process and accelerates the prefill time. Locret is compatible with all decoder-only LLMs and imposes minimal additional hardware requirements.
Contribution 3: We extensively evaluate Locret, demonstrating its ability to achieve comparable performance with full KV cache while maintaining inference efficiency. Locret achieves over 20×20$\times$20 × and 8×8$\times$8 × KV cache compression ratios for Phi-3-mini-128K and Llama-3.1- 8B-instruct, respectively. Additionally, Locret-Q accelerates prefill by over 2×2$\times$2 × on query-driven tasks without significant performance degradation. This framework enables full comprehension of long contexts on consumer-grade devices without compromising the generation quality, and introduces minimal additional system optimizations.
@@ 2 Related Work
This paper focuses on optimizing long-context LLM inference. Existing efforts can be categorized into algorithm and system optimization. For more details about LLMs, please refer to the surveys (Zhao et al., 2023; Lu et al., 2024).
@@
Algorithm Optimizations aim to reduce the size of the KV cache and can generally be classified into three categories: quantization-based methods, sparsity-based methods, and token dropping methods. Quantization-based methods (Liu et al., 2024b; Hooper et al., 2024; Zandieh et al., 2024; Zhang et al., 2024a) use low-bit values to represent the KV cache, reducing cache memory overhead and improving cache computing efficiency. These quantization-based methods suffer from hardware-oriented operator customization and additional inverse quantization overhead. Sparsity-based methods (Ge et al., 2024; Jiang et al., 2024a; Yang et al., 2024; Lou et al., 2024; Lv et al., 2024) leverage the sparsity patterns of attention heads to reduce both computational and I/O costs. Combining different patterns can yield further optimization by identifying specific patterns for each head (Ge et al., 2024; Jiang et al., 2024a; Xiao et al., 2024b). For more details on sparsity-based methods, please refer to the surveys (Yuan et al., 2024; Kang et al., 2024; Shi et al., 2024). Although quantization-based methods and sparsity-based methods have achieved promising results, they cannot address the issue that the KV cache memory overhead increases linearly with the context length. Eviction-based methods, such as H2O (Zhang et al., 2024d), ScissorHands (Liu et al., 2024a), and SirLLM (Yao et al., 2024), rank KV cache units by certain statistical metrics to identify the most influential units, discarding others to reduce memory usage and speed up attention computation. Pooling-based methods (Nawrot et al., 2024; Rajput et al., 2024), especially StreamingLLM (Xiao et al., 2024c) and LoCoCo (Cai et al., 2024a), compress multiple adjacent KV cache units into a single unit using pre-designed transformations. More important units will merge into compressed units with higher weights. Eviction-based and pooling-based methods drop or merge tokens to maintain a static cache size, but struggle with accurate victim selection and optimal pooling function design.
@@
System Optimizations alleviate the challenge of long-context inference from a system-level perspective, by fully considering hardware features. Offloading-based methods (Sheng et al., 2023; Xiao et al., 2024a; Wu et al., 2024; Sun et al., 2024) use CPU memory to store the KV cache and retrieve only the most relevant chunks to GPU memory before inferring a new chunk. These methods reduce maximum GPU memory usage at the cost of introducing CPU-GPU communication overhead. Hardware-aware methods, such as Flash-Atttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and Page-Attention (Kwon et al., 2023), enable more efficient runtime memory management by considering GPU architectures (Ghorpade et al., 2012). In addition, building inference infrastructures with a more efficient programming language (llama.cpp, ; llama2.c, ; rustformers, ), or adopting disaggregated inference (Jiang et al., 2024b; Zhong et al., 2024; Qin et al., 2024a; Hu et al., 2024a), can also greatly improve long-context inference efficiency. Since system optimizations primarily enhance efficiency by leveraging hardware resources rather than directly optimizing attention patterns or the KV cache, relying solely on them cannot adequately address the challenges of long-context LLM inference. Several efforts have integrated algorithm optimizations into system optimizations (Agrawal et al., 2023; Lee et al., 2024), such as KTransformers (KVCache.AI, 2024) leveraging offloading based on InfLLM (Xiao et al., 2024a), achieving promising results.
Figure 2: The framework of Locret. “$\mathbf{R}$” represents the retaining head. $\text{P}_{i}$ and $\text{A}_{i}$ correspond to the $i$-th prompt token and answer token. “t” represents the time step in chunked prefill, “$b$” represents the budget size, and “$n_{s}$” represents the length of the stabilizers. For simplicity, our notation here does not reflect the concept of layers.
3 Methodology of Locret
@@ 3.1 Preliminaries
Transformer Architecture. Given a token sequence $\{t_{1},\cdots,t_{n}\}$ as the input prompt of the transformer-based LLM, we denote the output hidden states of the $i$-th layer as $\textbf{H}^{(i)}$, and denote the input embeddings of the first layer as $\textbf{H}^{(0)}$. For each transformer layer, it consists of an attention block and a feedforward neural network (FFN) block. Attention blocks often follow the grouped-query attention (GQA) architecture (Ainslie et al., 2023), with $h$ query heads and $h/g$ KV heads, where $g$ is the group size, i.e., $g$ heads share the same KV heads. The multi-head attention (MHA) architecture adopted in the original transformer can be regarded as a special GQA ($g=1$). In the $i$-th layer, the attention score of the $j$-th query head is formalized as $\mathbf{A}_{j}^{(i)}=\texttt{softmax}\left(\mathbf{Q}_{j}^{(i)}{\mathbf{K}_{\lceil{j/g}\rceil}^{(i)\top}}/\sqrt{d_{m}}\right)\cdot\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}$ , where $d_{m}$ represents the hidden size for each head and $\left[\mathbf{Q}_{j}^{(i)},\mathbf{K}_{\lceil{j/g}\rceil}^{(i)},\mathbf{V}_{\lceil{j/g}\rceil}^{(i)}\right]=\mathbf{H}^{(i-1)}\cdot\left[{\mathbf{W}_{j}^{(i)\texttt{Q}}},{\mathbf{W}_{\lceil{j/g}\rceil}^{(i)\texttt{K}}},{\mathbf{W}_{\lceil{j/g}\rceil}^{(i)\texttt{V}}}\right]$ After obtaining the attention score, the output of the $i$-th attention block is $\mathbf{A}^{(i)}=\left[\mathbf{A}_{1}^{(i)},\cdots,\mathbf{A}_{h}^{(i)}\right]\cdot\mathbf{W}^{(i)\texttt{O}}$ , and the output hidden states of the $i$-th layer is $\mathbf{H}^{(i)}=\texttt{FFN}(\mathbf{A}^{(i)})$
@@
KV Cache and Chunked Prefill. Given the input prompt sequence $\{t_{1},\cdots,t_{n}\}$, during the prefill stage, all prompt tokens are processed in a single forward pass. After the prefill, $\mathbf{K}^{(i)}=\left[\mathbf{K}^{(i)}_{1},\cdots,\mathbf{K}^{(i)}_{h/g}\right]$ and $\mathbf{Q}^{(i)}=\left[\mathbf{Q}^{(i)}_{1},\cdots,\mathbf{Q}^{(i)}_{h/g}\right]$ are stored as the KV cache, whose sequence length is $n$. During the decoding stage, each time a token is decoded, a forward pass is conducted only for this token and decode the next token. In this process, the KV cache is used to avoid redundant attention computation. Chunked prefill is a method for reducing peak memory usage by spliting tokens into mulitple chunks and prefilling tokens chunk by chunk. Taking both the KV cache and chunked prefill into account, the attention block can be modified as follows,
$(\displaystyle\mathbf{A}{[n+1\colon n+B]}_{j}^{(i)}=\texttt{softmax}\Biggl{(}$1)
$\frac{\mathbf{Q}{[n+1\colon n+B]}_{j}^{(i)}\mathbf{K}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)\top}}{{\sqrt{d_{m}}}}\Biggr{)}$ $\mathbf{V}{[1\colon n+B]}_{\lceil{j/g}\rceil}^{(i)},$
where $\mathbf{A}[n+1\colon n+B]$ denotes the attention output for the tokens $\{n+1,\cdots,n+B\}$, and $B$ is the number of tokens processed in a single forward pass. For decoding, $B=1$, while for chunked prefill, $B$ corresponds to the chunk size. For the $k$-th token in the context, its attention output is $\mathbf{A}[k]$, its key and value vectors are $\mathbf{K}[k]$ and $\mathbf{V}[k]$.
@@
Cache Eviction. In the cache eviction process, we treat the KV vector pair of a single token within one attention head as the smallest cache unit. We denote the cache unit of the $k$-th token as $c_{k}=(\mathbf{K}[k],\mathbf{V}[k])$. Assuming a memory budget $b$, representing the maximum number of cache units that can be stored, the abstract form of the attention block can be written as $c_{k}=f(c_{1},c_{2},\cdots,c_{k-1})$. With limited cache capacity, this process can only be approximated by $\tilde{c}_{k}=f(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b^{\prime}\leq b$, and $p_{1},p_{2}\cdots,p_{b^{\prime}}\in\{1,2,\cdots,k-1\}$. When the cache is full, one unit must be evicted. We select the victim using some policy $p_{v}=\text{Policy}(\tilde{c}_{p_{1}},\cdots,\tilde{c}_{p_{b}};\tilde{c}_{k})$, and the key challenge is to develop an effective policy that minimizes the error $\|\tilde{c}_{k}-c_{k}\|$.
3.2 Framework of Locret